40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Entry  Fri May 26 21:53:20 2017, Koji, General, Configuration, Trans RF PD setup 
    Reply  Tue Jun 6 00:49:48 2017, Koji, General, Configuration, Trans RF PD setup 
       Reply  Tue Jun 6 22:00:36 2017, Koji, General, Configuration, Trans RF PD setup 
Message ID: 279     Entry time: Tue Jun 6 00:49:48 2017     In reply to: 278     Reply to this: 280
Author: Koji 
Type: General 
Category: Configuration 
Subject: Trans RF PD setup 

Last week, I further worked on the RF system to install 20dB coupler on the agilent unit and setup the R channel. This allowed me to make the FSR/TMS measurement of the OMC.

And today several optical improvement has been done.

- The input/output fiber couplers were adjusted to have the maximum transmission through the PBS right before the OMC.
- The HWP on the output side of the faraday was adjusted to have ~40mW input to the OMC.

Then, the OMC curved mirror is now held by the new in-situ gluing fixture instead of the conventional fixture attached upside down.
The OMC was ocked again and the input alignment was adjutsed. The fixture is blocking the QPD path, so it's not possible to confirm the proper alignment of the cavity (w.r.t. the QPD paths).

The precise positions of the spots could not be confirmed as the battery of the IR viewer was empty. Quick check of the spots by the card tells that the spot on the CM2 (PD side) is slightly too close to FM2 (output coupler). I wonder if this could be solved by rotating the curved mirror.

Otherwise everything look good. Let's try to glue the curved mirror tomorrow.

Note: Spot on CM2 is too close to the edge of the hole on the mounting prism. The meausrementof CM1 is telling that the curverture center is located 2.7mm upper side of the center of the mirror if the HR side arrow is up (and it is the case). If we move the arrow to the QPD path side (90deg CW viewed from the face side), this corresponds to ~1.1mrad CCW tilt in Yaw (viewed from the top of the prism). According to the matrix calculation (T1500060) this will induce ~1.5mm shift of the beam. This should be tried before gluing.

ELOG V3.1.3-